OpenStudy (anonymous):

A store sells yellow and white hats. The white hats cost $8 more than the yellow ones. In one day the store sold 5 yellow hats and 8 white ones for a total income of $207. How much did each color of hats cost?

OpenStudy (anonymous):

ok, this is a system of equations you first have to make an equation for the price for the white hats, W= Y+8, then an eqn for the total price, 5Y+8W=207, then you subsitute the W eqn into the total price eqn, 5Y+8(Y+8)=207 5Y+8Y+64=207 13Y+64=207 (then subtract 64 on both sides) 13Y=143 Y=11, the price of yellow hats is $11 go back to the W eqn and plug in 11 for Y and add 8, so the price of white hates is $19

OpenStudy (anonymous):

*hats

OpenStudy (anonymous):

Amazing! Thanks!

OpenStudy (anonymous):

np :)